OpenStudy (anonymous):

Simplify x^2-12x+20/x-10: A. x-2 B. x-2/x-10 C. x-2/10 D. x-10/x-2

OpenStudy (anonymous):

Is that \[ \frac{x^2 - 12x + 20}{x - 10}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[=((x-10)(x-2))/(x-10)\] when\[x \neq 10\] =\[x-2\]

OpenStudy (anonymous):

So what you have up the top is a quadratic. We can factorise this and see what we get. \[\frac{(x - 10)(x - 2)}{x - 10}\] Can you see any way to simplify this?

OpenStudy (anonymous):

Cancel (x-10) on both sides to leave the answer as (x-2)

OpenStudy (anonymous):

Yep. Well done.

OpenStudy (anonymous):

Thank you!